OpenStudy (anonymous):

find the slope-intercept form for the line satisfying the following x-intercept 2, y-intercept 2/5

OpenStudy (jim766):

the x intercept is the point where y = 0 the y intercept is the point where x = 0 can you write the two points?

OpenStudy (anonymous):

(2,0),(0,2/5)

OpenStudy (jim766):

right, not to write the eq of line, you need a slope and a y intercept they already gave you the y intercept all you need is the slope. can you find it given the two points?

OpenStudy (jim766):

y2-y1/x2-x1

OpenStudy (anonymous):

y=1x+0

OpenStudy (jim766):

(2/5) - 0/ 0 - 2 = -1/5

OpenStudy (jim766):

-1/5 is the slope , the y intercept is 2/5, just substitute into y = mx + b

OpenStudy (anonymous):

i don't get it

OpenStudy (jim766):

y = -1/5x + 2/5

OpenStudy (jim766):

what part dont you get?

OpenStudy (anonymous):

it just seemed more difficult than it really is.

OpenStudy (anonymous):

thanks